What mass of oxygen is needed for the complete combustion of 4.00×10−3 g of methane?
1
Expert's answer
2016-09-16T08:59:04-0400
The combustion reaction of methane:
CH4 + 2O2 → CO2 + 2H2O
1 mol 2 mol 1 mol 2 mol
The number of moles CH4:
n(CH4) = m(CH4) / M(CH4) = 4.00×10−3 / 16 g/mol = 0.00025 mol
The number of moles O2:
n(O2) = n(CH4) × 2 = 0.00025 mol × 2 = 0.0005 mol
m (O2) = n(O2) × M (O2) = 0.0005 mol × 32 g/mol = 0.016 g
Answer: 0.016 g